OpenStudy (anonymous):

how to make -3√(x-6)+ 7=14? Giving a gold medal

OpenStudy (anonymous):

First thing you have to do is to transpose 7 to the right side or subtract both sides by 7. can you do that?

OpenStudy (anonymous):

Yeah,

OpenStudy (anonymous):

what will you get?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

you'll end up with something like this: \(-3 \sqrt{(x-6)}=14-7\\-3 \sqrt{(x-6)}=7 \) now divide both sides by -3

OpenStudy (anonymous):

radical(x - 6) = 2.repeating3

OpenStudy (anonymous):

How do I give u a medal?

OpenStudy (anonymous):

use the equation or draw tool, so that you can post your solution properly

OpenStudy (mathstudent55):

@bigjimneedshelp Remember that positive divided by negative is negative, so be careful with 7 divided by -3.